-3z/4=9/16
I have the answer -3/4
I want to know how to find the answer

Answer by richwmiller(17219)

multiply both sides by 16
-12z=9
divide both sides by -12
z=9/-12
reduce
z=-3*3/3*4
z=3/3*-3/4
z=-3/4